OpenStudy (anonymous):

Find f'(x) and f'(c). MEDAL WILL BE REWARDED!!! Function: f(x)=(x^2-2x+1)(x^3-1) Value of c: 1

OpenStudy (anonymous):

f(x)=(x^2-2x+1)(x^3-1) f(x)=x^5 -2x^4 +x^3 -x^2 +2x -1 from here you can differentiate the function :) do yyou think you can do it?

OpenStudy (anonymous):

well you could of also not of expanded it and just used product rule but i think this is easier

OpenStudy (anonymous):

yes thank you! Ill tag you in a question if I have any problems!

OpenStudy (anonymous):

np!

OpenStudy (anonymous):

@HawkCrimson Is the final answer f'(x)=5x^4-8x^3+3x^2-2x+2?

OpenStudy (anonymous):

yep! now the last part :)

OpenStudy (anonymous):

Whats the last part? @HawkCrimson